4(7s-3)-7(3s-5)= 0
how do you solve?

Answer by checkley77(12844)

4(7s-3)-7(3s-5)=0
28s-12-21s+35=0
7s=12-35
7s=-23
s=-23/7 ans.
Proof:
4(7*-23/7-3)-7(3*-23/7-5)=0
4(-161/7-3)-7(-69/7-5)=0
4(-161-3*7)/7-7(-69-5*7)/7=0
4(-161-21)/7-7(-69-35)/7=0
4(-182/7)-(-104)=0
-728/7+104=0
-104+104=0
0=0